OpenStudy (anonymous):

Physics Kirchhoff's rules; finding voltage in series circuit given the electrical potential difference at battery with loads with unknown resistance. Thanks

OpenStudy (anonymous):

Based on the diagram below

OpenStudy (anonymous):

I need to calculate the voltage at Delta V given the electrical potential differences necessitated at each load station to make a circuit. I know that 60V will be the starting point but given the "unknown" resistance I am slightly void of reason

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (rajat97):

what does the statement of Kirchoff's loop rule say??

OpenStudy (anonymous):

I am aware that principle of resistance in series circuit assumes Rt=R1+R2+R3...

OpenStudy (anonymous):

Right? @ganeshie8

OpenStudy (rajat97):

i think that there is no need to use that equation you can do it by only using kirchoff's loop law it says that if you loop the circuit from a starting point and then come to that point again, the total potential fall or rise should be 0 Volts

OpenStudy (rajat97):

or mathematically, the sum of all the potential differences accross each device (resistance,battery,etc.) should be 0 did you get it now??

OpenStudy (rajat97):

thanks for the medal @Vocaloid

OpenStudy (anonymous):

Really? Adding them all makes it all zero? However I don't see how voltage will be summed to zero given all of them have positive values.

OpenStudy (anonymous):

So from your reasoning v2=60-1.1-2.5=56.4

OpenStudy (anonymous):

That's a simple Kingdom Come

OpenStudy (anonymous):

Or is it really?

OpenStudy (rajat97):

yes that's what i mean!

OpenStudy (rajat97):

i'm sorry it took too long to reply

OpenStudy (anonymous):

Ok @rajat97 If I've seen further it is by standing on the shoulders of giants like you

OpenStudy (anonymous):

XD LOL

OpenStudy (rajat97):

i didn't get that i mean what did you mean to say? :)

OpenStudy (anonymous):

Do you realize we are almost equal to or exceeding the conceptual understanding of Kirchoff right now at the time? That means technically we are more informed than our past time physicistsXD

OpenStudy (anonymous):

world is getting smarter ever more

OpenStudy (rajat97):

hahahaha but to formulate the law at that time was not an easy nut to crack:)

OpenStudy (anonymous):